Recent Log Entries

PD UPDATE: beam trips

Log Number: 3847817Logbook(s): Tags: Entry Makers: beneschBody: 

Beam trips for all halls were greatly elevated by an unapproved Hall D raster test. No ATLis was submitted so it was not scheduled or made known to the PD. It has been terminated.

Active collimator testing (Part 2)

Log Number: 3847802Logbook(s): Entry Makers: michaelmBody: 

Fig. 1-4: 6.2A x2.2A on BS04H/V back on the diamond
Fig. 5/6: AC X target: -0.1 > 1.0 (substantial peak IC rate reduction)... can get to about 6.4A (H) by 2.2A (V) before encountering trips. Moving to larger X is seemingly not effective.

Update: PD called in an requested the test be terminated as an ATLis task is not in place. Beam has terminated and rastering turned off. Backing all lock targets and positions back out to initial configurations.

9/14/20 Hall D raster testing - setup (Part 1)

Log Number: 3847778Logbook(s): Entry Makers: michaelmBody: 

Under direction of Richard Jones:
Fig. 1-3: Rastering off.
Fig. 4/5: Rastering on: BS04H(23Hz & 6.2A) / BS04V(27 Hz & 2.2 A) we are limited by the present diamond ion chamber settings. Modest steering in the line brought slightly more range, but I think we are running into the diamond frame a bit. Hall D now moving to an amorphous radiator.
Fig. 6-8: Now on 4.5e-4 Al - raster off.
Fig. 9-12: Raster on at previouse settings, rates much better.
Fig. 13-15: Square rastering on Al. 7.8 A (H) x 2.2 A (Y) produces a ~6x6mm box at IPM5C11B.

Downtime Incident CLOSED: Hall A Access

Log Number: 3847726Logbook(s): Body: IDDescriptionPeriodCause8698Title:
Hall A Access

Summary:
Hall A Access; sieve slit repositioningDuration:
17 minutes

Time Down:
14-Sep-2020 02:03

Time Up:
14-Sep-2020 02:20

Pages